I need help with this...
Write the slope-intercept form of an equation of the line that passes through thr given point and is perpendicular to the graph of each equation.
(-2,-2), y= -1/3x+9
I got really confused in class today and my teacher explained how to do these problems in examples. Is there a formula i can use for it?

Found 2 solutions by 303795, checkley71:
Answer by 303795(602)

y= -1/3x+9 The slope of this equation is -1/3.
The slopes of two lines which are perpendicular multiply to give -1 so the slope of the line perpendicular to this one multiplies by -1/3 to give -1. The new slope will be 3 because 3 x -1/3 = -1.
The new line passes through the point (-2,-2) so this gives you an x value (-2) and a y value (-2).
The form of the line is
y=mx + b (or some people use c)
The only value not known in the equation is the b so
-2=3*(-2)+b
-2=-6 + b
b=-2+6=4
so the new line will be y = 3x + 4

Answer by checkley71(8403)

Y=-1/X+9 IS A LINE WITH A SLOPE OF (-1/3) & A Y INTERCEPT OF (0,9).
FOR A LINE TO BE PERPENDICULAR IT MUST HAVE A SLOPE THAT IS THE NEGATIVE RECIPRICAL OF THE LINE IT IS PERPENDICULAT TO. THUS THE SLOPE IS 3 FOR THE PERPENDICULAR LINE THROUGH (-2,-2).
NOW SUBSTITUTE THESE POINTS X=-2, Y=-2 AND THE SLOPE IN THE LINE FORMULA
Y=mX+b WE GET
-2=3(-2)+b
-2=-6+b
B=-2+6
b=4 SOLUTION FOR THE Y INTERCEPT (0,4)
THUS THE EQUATION FOR THIS LINE IS
Y=3X+4